OpenStudy (anonymous):

solve the system by the method of substitution 5x-y=5 5x-3y=15 Options A-(0,-5) B-(-5,0) C-(5,1) D-(1,5) Please show work thanks PLEASE HELP ME??????

OpenStudy (anonymous):

first solve equation 1 for y 5x-y=5 y=5x-5 is it ok till now

OpenStudy (anonymous):

the second one would be y=5/3x-5

OpenStudy (anonymous):

you do not need the second one now . just follow me is this ok with you.? first equation is 5x-y=5 y becomes y=5x-5 tell me .

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but i forget how do u change those types of problems the one that is 5x-y=5

OpenStudy (anonymous):

ok first equation is 5x-y=5 add y to both sides 5x-y+y=5+y 5x=5+y now subtract 5 from both sides 5x-5=5+y-5 5x-5=y y=5x-5 i hope it is clear now . now put the above value of y in the second equation at the place of y 5x-3y=15 since y=5x-5 5x-3(5x-5)=15 5x-15x+15=15 -10x=0 x=0/10 x=0 is it ok till now ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

put x=0 in the any of the given equation and tell me the value of y put in the first one x=0 5x-y=5 tell me y value.

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

so the answer is a

OpenStudy (anonymous):

yes .you are genius !! well done ..

OpenStudy (anonymous):

thnks i wouldnt be a genius if it wasnt for u

hero (hero):

I was just going to say why not just let 5x = 5x and set 5 + y = 15 + 3y

OpenStudy (anonymous):

wat

hero (hero):

Both equations have 5x, so you can isolate 5x in each

OpenStudy (anonymous):

yea u could do that

OpenStudy (anonymous):

yes @Hero you are right. there always more than one way to solve the question. you can subtract the two equations as well . at least right now i know more than seven methods by which it can solved !! :)

hero (hero):

Yes, that's great.